my question is: equation of line that contains (-6,1) and is perpendicular to the graph of y=1/3x+8
Answer by robertb(5830)

The slope of the perpendicular line has to be the negative reciprocal of 1/3 (the slope of the given line), and so is -3. The line then has form y = -3x + b. To solve for b, use the coordinates of the given point: 1 = -3*-6 + b, or 1 = 18 + b, or b = -17. The equation is then y = -3x-17.